P5-3A Presented here are selected transactions for Norlan Inc. during September of the current year. Norlan uses a perpetual inventory system. Sept. 2 Purchased equipment on account for $65,000, terms n/30, FOB destination. 3 Freight charges of $950 were paid by the appropriate party on the September 2 purchase of equipment. 4 Purchased supplies for $4,000 cash. 6 Purchased inventory on account from Hillary Corp. at a cost of $65,000, terms 1/15, n/30, FOB shipping point. 7 Freight charges of $1,600 were paid by the appropriate party on the September 6 inventory purchase. 8 Returned damaged goods costing $5,000 that were originally purchased from Hillary on September 6. Received a credit on account. 9 Sold goods costing $15,000 to Fischer Limited for $20,000 on account, terms 2/10, n/30, FOB destination. 10 Freight charges of $375 were paid by the appropriate party on the September 9 sale of inventory. 17 Received the balance due from Fischer. 20 Paid Hillary the balance due. 21 Purchased inventory for $6,000 cash. 22 Sold inventory costing $20,000 to Kun-Tai Inc. for $27,000 on account, terms n/30, FOB shipping point. 23 Freight charges of $500 were paid by the appropriate party on the September 22 sale of inventory. 28 Kun-Tai returned goods sold for $10,000 that cost $7,500. The merchandise was restored to inventory. Instructions (a) Record the September transactions on Norlan’s books. (b) Assume that Norlan did not take advantage of the 1% purchase discount offered by Hillary Corp. and paid Hillary on October 3 instead of September 20. Record the entry that Norlan would make on October 3 and determine the cost of missing this purchase discount to Norlan.  

P5-12A Data for Norlan Inc. are presented in P5–3A. Instructions (a) Record the September transactions on Norlan’s books, assuming it uses a periodic inventory system instead of a perpetual inventory system. (b) Assume that Norlan did not take advantage of the 1% purchase discount offered by Hillary Corp. and paid Hillary on October 3 instead of September 20. Record the entry that Norlan would make on October 3 and determine the cost of missing this purchase discount to Norlan.
Answer #1

P5-3A

Date Account Titles Debit Credit
Sep-02 Equipment $                65,000
      Accounts Payable $                65,000
(Equipment purchased on account)
Sep-03 Equipment $                      950
      Cash $                      950
(Payment of freight on equipment)
Sep-04 Supplies $                  4,000
      Cash $                  4,000
(Supplies purchased for cash)
Sep-06 Inventory $                65,000
     Accounts Payable $                65,000
(Payment of inventory on account)
Sep-07 Inventory $                  1,600
     Cash $                  1,600
(Paid freight in cash)
Sep-08 Accounts Payable $                  5,000
      Inventory $                  5,000
(Returnd damaged goods)
Sep-09 Accounts Receivable $                20,000
      Sales Revenue $                20,000
(Sales made on account)
Cost of Goods Sold $                15,000
      Inventory $                15,000
(Cost of goods sold recorded)
Sep-10 Delivery Expense $                      375
      Cash $                      375
(Paid freight on FOB Destination sale)
Sep-17 Cash $                19,600 =20000*98%
Sales Discount $                      400 =20000*2%
       Accounts receivable $                20,000
(Cash collected net of 2% discount)
Sep-20 Accounts Payable $                60,000 =65000-5000
      Cash $                59,400 =60000*99%
      Inventory $                      600 =60000*1%
(Paid net of 1% discount)
Sep-21 Inventory $                  6,000
      Cash $                  6,000
(Purchased inventory for cash)
Sep-22 Accounts Receivable $                27,000
      Sales Revenue $                27,000
(Sales made on account)
Cost of Goods Sold $                20,000
      Inventory $                20,000
(Cost of goods sold recorded)
Sep-23 No entry
Sep-28 Sales Returns and allowances $                10,000
      Accounts receivable $                10,000
(Sales returns)
Inventory $                  7,500
      Cost of Goods Sold $                  7,500
(Inventory added back)

Part 5-12A

Date Account Titles Debit Credit
Sep-02 Equipment $                65,000
      Accounts Payable $                65,000
(Equipment purchased on account)
Sep-03 Equipment $                      950
      Cash $                      950
(Payment of freight on equipment)
Sep-04 Supplies $                  4,000
      Cash $                  4,000
(Supplies purchased for cash)
Sep-06 Purchases $                65,000
     Accounts Payable $                65,000
(Payment of inventory on account)
Sep-07 Freight In $                  1,600
     Cash $                  1,600
(Paid freight in cash)
Sep-08 Accounts Payable $                  5,000
      Purchase Returns and allowances $                  5,000
(Returned damaged goods)
Sep-09 Accounts Receivable $                20,000
      Sales Revenue $                20,000
(Sales made on account)
Sep-10 Delivery Expense $                      375
      Cash $                      375
(Paid freight on FOB Destination sale)
Sep-17 Cash $                19,600 =20000*98%
Sales Discount $                      400 =20000*2%
       Accounts receivable $                20,000
(Cash collected net of 2% discount)
Sep-20 Accounts Payable $                60,000 =65000-5000
      Cash $                59,400 =60000*99%
      Purchase Discount $                      600 =60000*1%
(Paid net of 1% discount)
Sep-21 Purchases $                  6,000
      Cash $                  6,000
(Purchased inventory for cash)
Sep-22 Accounts Receivable $                27,000
      Sales Revenue $                27,000
(Sales made on account)
Sep-23 No entry
Sep-28 Sales Returns and allowances $                10,000
      Accounts receivable $                10,000
(Sales returns)
Date Account Titles Debit Credit
Oct-03 Accounts Payable $                60,000
      Cash $                60,000

Cost of missing purchase discount = $600
In percentage = 1%/99% x 365/13 = 28.36%
Since payment is made by 13 days late from end of discount period, denominator is taken as 13 days